For some reason I can not set up these kind of word problems:
According to data from the US dept of defense and congressional research service, the estimated annual cost y (in thousands of dollars) per active duty service member in the us armed forces in year x is given by the equation:
6000x-2000y+136000=0 (4< x < 18)
where x=4 corresponds to 1994. find the year in which the estimated annual cost per service member is (a) $92,000 (b) 110,000
-i have tried setting up the equation by substituting 4 for x and then entering both equations in my TI 83. I believe it's an intersection answer between two equations, however, i am not getting anywhere and i have attemped four other similar questions hoping i something would click. There are also no examples in the text i am using to assist. Thank you for any help you may provide.
erin mcavoy

Found 2 solutions by stanbon, xcentaur:
Answer by stanbon(48568)

Check your equation; it simplifies to y=3x+68 , which is hardly reasonable.
Cheers,
Stan H.

Answer by xcentaur(357)

6000x-2000y+136000=0 (4< x < 18)
where x=4 corresponds to 1994


It reduces to y=3x+68.


x=4,y=3(4)+68=12+68=80
x=5,y=3(5)+68=15+68=83
x=6,y=86
x=7,y=89
x=8,y=92------------->(a)
x=9,y=95
x=10,y=98
x=11,y=101
x=12,y=104
x=13,y=107
x=14,y=110----------->(b)
x=15,y=113
x=16,y=116
x=17,y=119
x=18,y=122


If you substitute x=0,we get y=68.Multiply this by 1000 and we get a credible figure,considering it as 1990. Then for x=18,y=122 ie in the year 2008, a figure of $122000 seems to make sense.


That would make 1998 and 2004 your answers,though of course we have no logical reasoning to back this up.


If this is a valid question,and you do manage to solve this,drop a mail at:
xcentaur[at]hotmail[dot]com
Hope this helps,
xC